### v2

@@ Abstract
Image restoration aims to recover a high-quality clean image from its degraded version. Recent progress in image restoration has demonstrated the effectiveness of All-in-One image restoration models in addressing various degradations simultaneously. However, these existing methods typically utilize the same parameters to tackle images with different degradation types, thus forcing the model to balance the performance between different tasks and limiting its performance on each task. To alleviate this issue, we propose HAIR, a Hypernetworks-based All-in-One Image Restoration method that dynamically generates parameters based on input images. Specifically, HAIR consists of two main components, i.e., Classifier and Hyper Selecting Net (HSN). The Classifier is a simple image classification network used to generate a Global Information Vector (GIV) that contains the degradation information of the input image, and the HSN is a simple fully-connected neural network that receives the GIV and outputs parameters for the corresponding modules. Extensive experiments demonstrate that HAIR can significantly improve the performance of existing image restoration models in a plug-and-play manner, both in single-task and all-in-one settings. Notably, our innovative model, Res-HAIR, which integrates HAIR into the well-known Restormer, can obtain superior or comparable performance compared with current state-of-the-art methods. Moreover, we theoretically demonstrate that our proposed HAIR requires fewer parameters in contrast to the prevalent All-in-One methodologies. The code is available at https://github.com/toummHus/HAIR.
@@ 1 Introduction
Image restoration stands as a pivotal challenge within the field of computer vision, dedicated to the reconstruction of high-quality images from their degraded states. The presence of adverse conditions such as noise, haze, or rain can severely diminish the effectiveness of images for a variety of applications, such as autonomous navigation [1, 2], augmented reality [3, 4, 5]. Therefore, developing robust image restoration methods is of great importance. The use of deep learning in this domain has made remarkable progress, as evidenced by a suite of recent methodologies [6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. Nonetheless, the predominant approach in current research is to employ task-specific models, each tailored to address a particular type of degradation [16, 7, 12, 13, 8]. This tailored approach, while precise, presents a constraint in terms of universality, as it restricts the applicability of models to scenarios with varied or unknown degradations [17, 18, 19]. To overcome this limitation, the research community has increasingly turned its attention to the development of All-in-One image restoration models. These models are designed to tackle various degradations using one single model. Pioneering efforts in this area [20, 21, 1, 22, 23, 24, 25, 26, 15] have utilized a variety of advanced techniques such as contrastive learning [20], meta-learning [25], visual prompting methods [24, 27, 28, 14]. These approaches have undoubtedly made substantial contributions to the field, propelling it forward with notable advancements.
@@
Figure 1: The difference between our method and previous methods. (a) Previous All-in-One image restoration methods. For images with various degradations, these methods utilize a single model with the same parameters. (b) Our proposed HAIR. Given a certain degraded image, we first use a fixed Extraction module to extract the features, and feed the extracted feature into the Hypernetwork to generate the dynamic parameters for the fusion module and finally get the restored image by the generated parameters.
@@
Despite the notable successes of these methods, they share a common drawback, namely, they rely on a single model with fixed parameters to address a variety of degradations. This one-size-fits-all approach can hinder the model’s effectiveness when dealing with multiple degradations simultaneously. For example, when viewed through the lens of frequency domain analysis, haze is characterized as low-frequency noise, in contrast to rain, which is considered high-frequency interference. An effective dehazing model acts as a low-pass filter, preserving high-frequency details, whereas deraining requires the opposite—enhancing the high-frequency components. Consequently, a model must strike a balance between these conflicting demands of different degradations, which may lead to suboptimal performance.
@@
To mitigate the aforementioned issue, we propose HAIR in this paper. Our fundamental concept is the dynamic generation of parameters (weights) based on the degradation information present in the input image. HAIR employs Hypernetworks [29], a trainable neural network that outputs parameters for other neural networks, to take the degradation information from the input image and produce corresponding parameters. Specifically, for a given unknown degraded image, we first utilize a Classifier, similar to those used in image classification networks, to obtain its Global Information Vector, which contains crucial discriminative information about various types of image degradation (as shown in Fig. 2). This vector is then passed to the Hyper Selecting Net, which is a Fully Connected Network, to generate the requisite parameters, as illustrated in Fig. 1. Therefore, our method can dynamically adapt to various degradations. With these dynamically parameterised modules, we ultimately achieve the restored image. In brief, our contributions include:
@@
  • We propose HAIR, a novel Hypernet-based All-in-One image restoration method capable of dynamically generating parameters based on the input image’s characteristics. HAIR consists of two components, i.e. Classifier and Hyper Selecting Nets, both of which function as a plug-in-and-play module. Extensive experiments demonstrate that HAIR can significantly improve the performance of existing image restoration models in a plug-and-play manner, both in single-task and all-in-one settings.
@@
  • We theoretically prove that, for a given small error threshold $\epsilon$ in image restoration tasks, HAIR requires fewer parameters compared to mainstream embedding based All-in-One methods like [20, 24, 26].
Figure 2: The figure presents tSNE plots of the degradation embeddings utilized in PromptIR [24] and our HAIR. Each distinct color represents a unique degradation type. It becomes evident that HAIR excels not only in recognizing various degradation types, such as noise, rain, and haze, but also in distinguishing between the same type of degradation at varying intensities, exemplified by noise with standard deviations of [15, 25, 50]. This capability underscores HAIR’s remarkable adaptability and generalization when confronted with diverse degradation scenarios.
2 Related Preliminaries
@@ 2.1 All-in-One Image Restoration
While single degradation methods do achieve great success [9, 10, 11, 7, 12, 13], blind all-in-one image restoration, which aims to utilize a single deep blind restoration model to tackle multiple types of degradation simultaneously without the prior information of the degradation of the input image, has gained more attention recently  [1, 2, 30, 24, 25, 26]. The seminal work AirNet [20] achieves blind All-in-One image restoration by employing contrastive learning to extract degradation representations from corrupted images, which are subsequently utilized to restore the clean image. Next, IDR [25] decomposes image degradations into their underlying physical principles, achieving All-in-One image restoration through a two-stage process based on meta-learning. With the rise of LLM [31], prompt-based learning [24, 27, 14] has also emerged as a promising direction in computer vision tasks. Despite the success these methods have achieved, they typically use the same parameters for distinct degradations and inject the degradation information as conditional embeddings into the model, which may impair the performance of the models. In contrast, our proposed HAIR can adaptively generate specific parameters for the given degraded image, resulting in better generalization to unknown degradations. [32] theoretically explains why using data-conditional Hypernetworks to generate adaptive parameters works better than injecting conditional embeddings into the model.
2.2 Image Classification and Feature Extraction
Image classification is a fundamental task in Computer Vision, with the goal of categorizing images based on their content [33]. Typically, image classification contains two key steps: feature extraction and classification. The primary challenge in this task is feature extraction, which involves converting an image into a vector that encapsulates its most critical information. While traditional methods relied on handcrafted techniques [34], the advent of neural networks has revolutionized feature extraction [35, 36, 37, 38]. Given that these vectors are capable of classifying hundreds of categories [39], they must contain the essential information of the images. This concept also extends to the classification of degraded images. For instance, if we categorize degraded images into different types, e.g. noisy, rainy, or hazy, and by levels of degradation like slightly noisy, moderately noisy, and very noisy, the feature vectors should be effective in classifying these variations. The rationale is that recognizing different degradations is generally less complex than identifying hundreds of distinct objects. Therefore, our proposed method employs a Classifier to obtain a Global Information Vector (GIV), that captures the rich global information of the image, including its degradation characteristics. (As shown in Fig. 2) This GIV is then utilized to generate parameters tailored to the specific degradation present in the image.
@@ 2.3 Data-conditioned Hypernetworks
Hypernetworks [29] are a class of neural networks designed to generate weight values (parameters) for other networks. They can be broadly classified into three main types: Task-conditioned, Data-conditioned, and Noise-conditioned hypernetworks [40]. Among these methods, Data-conditioned hypernetworks are particularly noteworthy for their ability to generate weights contingent upon the distinctive features of the input data. This capability allows the neural network to dynamically adapt its behaviour in response to specific input patterns or characteristics, fostering flexibility and adaptability within the model. Consequently, this results in enhanced generalization and robustness. Data-conditioned Hypernetworks have found applications across the field of computer vision, including semantic segmentation [41] and image editing [42]. Despite previous attempts to integrate Hypernetworks into Image Restoration [43], these have primarily leveraged Task-conditioned hypernetworks. Such an approach cannot fully harness the potential of hypernetworks for the task of image restoration. To the best of our knowledge, our work first introduces Data-conditioned hypernetworks to the domain of image restoration, unlocking new possibilities for the field.
Figure 3: The overall framework of Res-HAIR. In this paper, we mainly incorporate HAIR into Restormer [12], which has demonstrated robust performance in all-in-one restoration scenarios [24, 26, 28]. HAIR introduces a Classifier that yields a Global Information Vector (GIV) $\mathbf{V_{g}}$ from the high-level feature $\mathbf{F_{g}}$ encapsulating the input image’s degradation information. Additionally, the Hyper Selecting Nets (HSNs) dynamically generate weights for the Transformer Blocks based on GIV, customizing the restoration process to address the specific degradation characteristics of each image. This innovative approach allows models to adapt to various image restoration tasks without altering the original architecture. For other baseline models, we follow a similar integration approach with HAIR: we introduce a classifier at the network’s midpoint and utilize the resulting GIV to dynamically generate weights for the subsequent layers of the network.
@@ 3 Method
In this section, we outline the core principles of our proposed effective all-in-one image restoration method. Note that we use Res-HAIR stands for Restormer embedded with HAIR.
Overall Pipeline As shown in Fig. 3, our network architecture is consistent with Restormer [12]. Given a certain degraded image input $\mathbf{X}\in\mathbb{R}^{H\times W\times 3}$, Res-HAIR utilizes a $3\times 3$ convolution to transform $\mathbf{X}$ into feature embeddings $\mathbf{F_{0}}\in\mathbb{R}^{H\times W\times C}$, where $C$ denotes the number of channels. These feature embeddings are then proceeded through a 4-level hierarchical encoder-decoder, resulting in deep features $\mathbf{F_{r}}$. Each level of the encoder-decoder incorporates several Transformer blocks, with an increasing number from the top to the bottom level, ensuring computational efficiency. The left three blue "Transformer Blocks" function as an Encoder, designed to extract features from $\mathbf{F_{0}}$ and ultimately produce a global feature map $\mathbf{F_{g}}\in\mathbb{R}^{\frac{W}{8}\times\frac{H}{8}\times 8C}$ with a large receptive field. $\mathbf{F_{g}}$ is then input into a Classifier to yield a Global Information Vector (GIV) $\mathbf{V_{g}}\in\mathbb{R}^{2C}$. The right four orange "HyperTrans Blocks" operate as a Decoder, aiming to adaptively fuse features at each Decoder level based on degradation, culminating in a restored image $\hat{\mathbf{X}}$. Specifically, at each Decoder level, the HyperTrans Block accepts two inputs: the input feature and the GIV $\mathbf{V_{g}}$. $\mathbf{V_{g}}$ is fed into the Hyper Selecting Net to generate weights for the corresponding Transformer Blocks, which are then applied to the input feature to produce the output. It is important to note that all HyperTrans Blocks use the same $\mathbf{V_{g}}$ derived from $\mathbf{F_{g}}$. Since the weights in the Decoder are generated based on $\mathbf{V_{g}}$, which contains the degradation information of $\mathbf{X}$, our method exhibits strong generalization and adaptability, leading to excellent restoration performance.
3.1 Classifier
As outlined in Section 2.2, our approach to extracting degradation information involves designing a straightforward Classifier akin to those used in image classification tasks. We start by entering $\mathbf{F_{g}}\in\mathbb{R}^{\frac{H}{8}\times\frac{W}{8}\times 8C}$ into the backbone, which is composed of multiple ResNet blocks [35] followed by downsampling ×$\times 2$. This process is aimed at downsizing the spatial resolution while distilling the essential information, as depicted by the following equations:
@@ 3.1 Classifier
$\mathbf{F_{g}^{t}}$ $=\text{Downsampling}(\text{ResNetBlock}(\mathbf{F_{g}^{t-1}}))),\quad\text{for}\ t=1,2,3.$
After three iterations, we obtain $\mathbf{F_{g}^{3}}\in\mathbb{R}^{\frac{H}{64}\times\frac{W}{64}\times 2C}$, which serves as the input for generating the GIV $\mathbf{V_{g}}\in\mathbb{R}^{2C}$:
@@
where GAP denotes Global Average Pooling. The inclusion of GAP ensures that $\mathbf{V_{g}}$ remains a one-dimensional vector of consistent size, irrespective of the input image’s resolution. Unlike traditional image classification networks, there is no requirement for a Softmax layer here, as its application would confine the values of $\mathbf{V_{g}}$ within the [0,1] range, potentially restricting the parameter generation process. Moreover, the backbone can be substituted with other straightforward image classification networks such as VGG [44] and Inception [45], without any detrimental impact on performance, given that the classification of degradation types is a relatively simple task.
3.2 Hyper Selecting Net
Upon obtaining the Global Information Vector $\mathbf{V_{g}}$ from the Classifier, it is fed into our Data-conditioned Hypernetworks, known as Hyper Selecting Nets, to generate weights for the corresponding Transformer Blocks. Given $\mathbf{V_{g}}\in\mathbb{R}^{2C}$, a Selecting Vector (SV) $\mathbf{V_{s}}\in\mathbb{R}^{N}$ is initially computed as follows:
$\mathbf{V_{s}}=\sigma(\mathbf{FCNN}(\mathbf{V_{g}})).$ (3)
Here, $\sigma$ stands for a Softmax operation and $\mathbf{FCNN}$ denotes a simple Fully-Connected Neural Network. Subsequently, the parameters $\mathbf{w}$ are derived as

### v4

@@ Abstract
Image restoration aims to recover a high-quality clean image from its degraded version. Recent progress in image restoration has demonstrated the effectiveness of All-in-One image restoration models in addressing various unknown degradations simultaneously. However, these existing methods typically utilize the same parameters to tackle images with different types of degradation, forcing the model to balance the performance between different tasks and limiting its performance on each task. To alleviate this issue, we propose HAIR, a Hypernetworks-based All-in-One Image Restoration plug-and-play method that generates parameters based on the input image and thus makes the model to adapt to specific degradation dynamically. Specifically, HAIR consists of two main components, i.e., Degradation-Aware Classifier (DAC) and Hyper Selecting Net (HSN). DAC is a simple image classification network used to generate a Global Information Vector (GIV) that contains the degradation information of the input image, and the HSN is a simple fully-connected neural network that receives the GIV and outputs parameters for the corresponding modules. Extensive experiments demonstrate that HAIR can significantly improve the performance of existing image restoration models in a plug-and-play manner, both in single-task and All-in-One settings. Notably, our proposed model Res-HAIR, which integrates HAIR into the well-known Restormer, can obtain superior performance compared with State-Of-The-Art (SOTA) methods. Moreover, we theoretically demonstrate that to achieve a given small enough error, our proposed HAIR requires fewer parameters in contrast to mainstream embedding-based All-in-One methods.
@@ 1 Introduction
Figure 1: Comparisions between our method and previous methods in the inference stage. (a) Previous All-in-One image restoration methods. These methods utilize a single model with fixed parameters to tackle different degradations. (b) Our proposed HAIR. Given a certain degraded image, we use Hypernetworks to generate the dynamic parameters for the decoder, and finally obtain the restored image. Note that "dynamic" and "fixed" in this paper are specially for the main networks, as discussed in Sec. 8.1.
(a) AirNet
(b) PromptIR
(c) HAIR (ours)
(d) HAIR (ours)
Figure 2: Comparison of tSNE plots for the degradation embeddings between previous methods and our HAIR (i.e. the GIVs). Each distinct color represents a unique degradation type. As shown in (c), our HAIR excels not only in recognizing various degradation types, such as noise, rain, and haze, but also in distinguishing between the same type of degradation at varying intensities, e.g. noise with different standard deviations. Even when confronted with composite degradations not encountered during training, HAIR can also accurately discriminate them, i.e. the GIVs for these composite cases located midway between the GIVs of their constituent degradations, as illustrated in (d).
Image restoration is an important task in the field of computer vision, aiming to reconstructing high-quality images from their degraded states. The presence of adverse conditions such as noise, haze, or rain can severely diminish the effectiveness of images for a variety of applications, such as autonomous navigation [48, 8], augmented reality [17, 42, 11]. Therefore, developing robust image restoration methods is of great importance. The use of deep learning in this domain has made remarkable progress, as evidenced by a suite of recent methodologies [64, 29, 6, 45, 25, 67, 60, 7, 28, 56, 13]. Nonetheless, the predominant approach in current research is to employ task-specific models, each tailored to address a particular type of degradation [66, 29, 60, 7, 6]. This tailored approach, while precise, presents a constraint in terms of universality, as it restricts the applicability of models to scenarios with varied or unknown degradations [57, 58, 39]. To overcome this limitation, many researchers have focused on developing All-in-One image restoration models recently. These models are designed to tackle various degradations using one single model. Pioneering efforts in this area [27, 59, 48, 30, 14, 38, 62, 56, 13] have utilized a variety of advanced techniques such as contrastive learning [27], meta-learning [62], visual prompting methods [38, 50, 13, 28, 9]. These approaches have undoubtedly made substantial contributions to this field.
@@
However, these All-in-One models share a common drawback, i.e. they rely on a single model with fixed parameters to address various degradations. This one-size-fits-all method can hinder the model’s effectiveness when dealing with multiple degradations simultaneously. For example, when viewed through the lens of frequency domain analysis, haze is characterized as low-frequency noise, in contrast to rain, which is considered high-frequency interference. An effective dehazing model acts as a low-pass filter, preserving high-frequency details, whereas deraining requires the opposite—enhancing the high-frequency components. Consequently, a model must balance these conflicting demands of different degradations, thus limiting its performance on each task. Further details on this issue are provided in Sec. 7.1.
@@
To mitigate the aforementioned issue, we propose a Hypernetworks-based All-in-One Image Restoration method (HAIR) in this paper. The core idea of HAIR is to generate the weight parameters based on the input image, and thus can dynamically adapt to different degradation information. HAIR employs Hypernetworks [19], a trainable neural network, to take the degradation information from the input image and produce the corresponding parameters. Specifically, for a given unknown degraded image, we first utilize a Degradation-Aware Classifier, similar to those used in image classification networks, to obtain its Global Information Vector, which contains crucial discriminative information about various types of image degradation (as shown in Fig. 2 (c-d)). This vector is then used to generate the necessary parameters, as illustrated in Fig. 1. With these dynamically parameterized modules, we ultimately achieve the restored image. In brief, our contributions include:
@@
  • We propose HAIR, a novel Hypernetworks-based All-in-One image restoration method that is capable of dynamically generating parameters based on the degradation information of input image. HAIR consists of two components, i.e. Degradation Aware Classifier and Hyper Selecting Nets, both of which function as plug-in-and-play modules. Extensive experiments indicate that HAIR can obviously improve the performance of existing models in a plug-and-play manner, both in single-task and All-in-One settings.
@@
  • We theoretically prove that, for a given small enough error threshold $\epsilon$ in image restoration tasks, HAIR requires fewer parameters compared to mainstream embedding-based All-in-One methods like [27, 38, 9].
2 Related Works
@@ 2.1 All-in-One Image Restoration
While single degradation methods do achieve great success [45, 25, 67, 29, 60, 7], All-in-One image restoration, which aims to utilize a single deep restoration model to tackle multiple types of degradation simultaneously without prior information about the degradation of the input image, has gained more attention recently  [48, 8, 23, 38, 62, 56]. The pioneer work AirNet [27] achieves All-in-One image restoration using contrastive learning to extract degradation representations from corrupted images. IDR [62] decomposes image degradations into their underlying physical principles, achieving All-in-One image restoration through a two-stage process based on meta-learning. With the rise of LLM [68], prompt-based learning has also emerged as a promising direction in image restoration tasks [38, 50, 28]. They typically produce a prompt embedding for each input image based on their content, then inject this prompt into the model to restore the image, which is essentially a conditional embedding. Among them, some works like DA-CLIP [32] and InstructIR [9] insightfully leverage pre-trained large-scale text-vision models to produce the visual prompts. Next, some methods like DaAIR [56] and AMIR [54] utilize routing techniques, which leverage multiple experts within the model to handle different degradation types or tasks by directing the input data along specialized pathways, to achieve adaptive image restoration. However, the output of these multiple experts is also essentially a kind of conditional embedding. Despite the success these methods have achieved, most of them typically use the same parameters for distinct degradations and only inject the degradation information as conditional embeddings into the model, which forces the model to balance different degradations, and thus impair the performance of the models. Tian et al. [47] recently employed low-rank matrix decomposition for weight modulation, but their method relies on pre-trained, task-specific weights and prior degradation information, lacking the dynamic parameter generation capability and adaptibility. In contrast, our proposed HAIR can dynamically generate specific parameters for the given degraded image using a hypernetwork and thus making the model adapt to unknown degradations better.
@@ 2.2 Data-conditioned Hypernetworks
Hypernetworks [19] are a class of neural networks designed to generate weights (parameters) for other networks. They can be classified into three types, i.e task-conditioned, data-conditioned, and noise-conditioned hypernetworks [5]. Among these methods, data-conditioned hypernetworks are particularly noteworthy for their ability to generate weights contingent upon the distinctive features of the input data. This capability allows the network to dynamically adapt its behaviour to specific input patterns or characteristics, fostering flexibility and adaptability within the model. Consequently, this results in enhanced generalization and robustness. Data-conditioned hypernetworks have been applied in many computer vision tasks, e.g., semantic segmentation [37] and image editing [2]. Despite previous attempts to integrate hypernetworks into image restoration [1, 14], these have primarily leveraged task-conditioned hypernetworks, which take a task embedding (like an embedding of derain) as the input and output the weights. They have two main drawbacks: (1) They require prior knowledge of the input image’s degradation type. (2) They cannot dynamically generate weights based on image content. Although Klocek et al. [24] tries to use Data-conditional Hypernetworks in Super-Resolution, it direcly maps a coordinate to a pixel, which is unreliable. To the best of our knowledge, our work is the first to introduce data-conditioned hypernetworks into the domain of All-in-One image restoration.
Figure 3: The overall framework of our proposed Res-HAIR. Res-HAIR is built by integrating our HAIR into the popular Restormer [60]. HAIR contains two modules, i.e. Degradation-Aware Classifier (DAC) and Hyper Selecting Net (HSN). The DAC is used to yield a Global Information Vector (GIV) $\mathbf{V_{g}}$ from the high-level feature $\mathbf{F_{g}}$ containing the degradation information of the input image. HSN dynamically generates Transformer Block weights based on GIV, tailoring restoration to each image’s unique degradation.
@@ 3 Method
In this section, we first introduce our proposed hypernetwork-based All-in-One image restoration module (HAIR) in detail. Then, we propose our All-in-One method (Res-HAIR) by integrating HAIR into the popular Restormer [60]. Finally, we provide a brief theoretical analysis on the proposed HAIR.
3.1 Proposed HAIR Module
In this subsection, we introduce HAIR, a hypernetworks-based All-in-One image restoration module designed to dynamically generate parameters for an image-to-image network based on the input image’s features in a plug-and-play manner. HAIR comprises two main components: the Degradation-Aware Classifier (DAC) and the Hyper Selecting Net (HSN). Next, we will introduce the two components in detail.
3.1.1 Degradation-Aware Classifier (DAC)
As outlined in Sec. 1, our approach to extracting degradation information involves designing a simple DAC akin to those used in image classification tasks. Taking Fig. 3 as an example, we start by entering the feature $\mathbf{F_{g}}\in\mathbb{R}^{\frac{H}{8}\times\frac{W}{8}\times 8C}$ into the backbone, which contains multiple ResNet blocks [20] followed by downsampling ×$\times 2$. This process is aimed at downsizing the spatial resolution while distilling the essential information, shown as follows:
@@ 3.1.1 Degradation-Aware Classifier (DAC)
$\mathbf{F_{g}^{t}}=\text{Downsampling}(\text{ResNetBlock}(\mathbf{F_{g}^{t-1}}))),$
where $t=1,2,3$. After three iterations, we obtain $\mathbf{F_{g}^{3}}\in\mathbb{R}^{\frac{H}{64}\times\frac{W}{64}\times 2C}$, which serves as the input of DAC for generating the global information vector (GIV) $\mathbf{V_{g}}\in\mathbb{R}^{2C}$ that captures the degraded information. Specifically,
@@
where GAP denotes Global Average Pooling. The inclusion of GAP ensures that $\mathbf{V_{g}}$ remains a one-dimensional vector of consistent size, irrespective of the input image’s resolution. Unlike traditional image classification networks, there is no requirement for a Softmax layer here, as its application would confine the values of $\mathbf{V_{g}}$ within the range [0,1], potentially restricting the parameter generation process. Moreover, the backbone can be substituted with other image classification networks, e.g., VGG [43] and Inception [44], with little negative impact on performance since degradation-type classification (discrimination) is a relatively simple task.
3.1.2 Hyper Selecting Net (HSN)
After obtaining GIV $\mathbf{V_{g}}$ from the DAC, it is then fed into our data-conditioned hypernetworks, i.e. Hyper Selecting Net (HSN), to generate weights for the corresponding modules. Let’s still take Fig. 3 as an example. Given $\mathbf{V_{g}}\in\mathbb{R}^{2C}$, a Selecting Vector $\mathbf{V_{s}}\in\mathbb{R}^{N}$ is initially computed as:
$\mathbf{V_{s}}=\sigma(\mathbf{FCNN}(\mathbf{V_{g}})),$ (3)
where $\sigma$ is a Softmax operation and $\mathbf{FCNN}$ denotes a simple fully-connected neural network. Unlike $\mathbf{V_{g}}$, $\mathbf{V_{s}}$ will be used to directly generate the parameters, so we need a $\sigma$ to make sure $\mathbf{V_{s}}$ is positive and $\in[0,1]$, thus making parameter generation more stable. Subsequently, the parameters